OpenStudy (anonymous):

Choose the factored form of the given equation. 6x2 + 2x - 8 = 0

OpenStudy (anonymous):

i mean 6x^2 + 2x - 8 =0

OpenStudy (anonymous):

options: a. (6x- 8)(x+ 1) = 0 b. (6x+ 1)(x- 8) = 0 c. (6x+ 8)(x- 1) = 0